You decide to have hot dogs for dinner. In the grocery store, you find that buns come only in packages of 12, whereas the hot dogs come in packages of 8. How can your purchases lead to having no buns or hot dogs left over? (How does this problem fit in with chemical bonding and other parts of this chapter?)
Answer #1

Answer -

Given,

Number of buns in a pack = 12

Number of hot dogs in a pack = 8

Number of packs of buns and hot dogs to be purchased to be use fully = ?

For this problem we have to find the number of packs of each in which the number of hot dogs and bun get equal.

this can be solved using LCM(Lowest Common Multiple)

2 12, 8
2 6, 4
2 3,2
3 3,1
1,1

LCM = 2* 2*2*3 = 24

Now Divide the LCM by the number of items in each pack,

24/12 = 2

24/8 = 3

So, if 2 packs of Buns and 3 packs of Hot dogs are purchased. This problem can be solved.

In the chemical bonding and other part of chemistry , this problem fits.

Because when a compound is formed, it forms in a way that the miniumum number of atoms are used and should be in whole numbers only. So, to balance reactions, etc this type of problem need to solved in the same way i.e used in this question.